OpenStudy (anonymous):

Solve x^2 + 10x - 11 = 0 by completing the square Ok, so I need help with this, I have the answer and I know how to do it, I just want to check it.

OpenStudy (campbell_st):

ok... to you need to group a few things \[x^2 + 10x = 11\] to complete the square find half of 10 and then square it... that's the value that makes the perfect square then add it to both sides..

OpenStudy (anonymous):

Ok that's a bit different then how I learned it but I think it's pretty much the same thing.. I would do x^2 + 10x + 25=11+25 Then group it to where it's (X+5)^2=36 Then you would put the (x+5)^2 under a radical sign, and the 36 under the radical sign, and you can't forget the plus and minus sign

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

But I don't know what to do after

OpenStudy (campbell_st):

ok... so when you take the square root there are 2 solutions, the positive and negative so take the square root of both sides \[x + 5 = \pm \sqrt{36} ~~or~~~x + 5 = \pm 6\] now subtract 5 from both sides of the equation and you get \[x = -5 \pm 6\] so you ahve 2 equations x = -5 -6 and x = -5 + 6 does that make sense

OpenStudy (anonymous):

Yes, see, I did all that But I had three instead of six. But thank you so much! I'm glad I checked with someone :)